# Visual Affordances: Enabling Robots to Understand Object Functionality

### v1

@@ VI-A Reproducibility challenges
TABLE X: Affordance Sheet. A tool inspired by Model Cards [30] to favour transparency and reproducibility of works for visual affordance predictions conditioned on robotic tasks.
MODEL NAME
Affordance task addressed OBJL FUNC FUNS EPE EIS
Datasets (RC1) Record link (where the datasets/benchmarks are stored)*
Licence: add licence here
Proposed method (RC2, RC3) Record link (where the weights of the model are stored)*
Code link (where model architecture and scripts are stored)
Model card [30]:
Licence: add licence here
Experimental setup (RC4) Data splits: add splits criteria here
Hyperparameters: add hyperparameters list here
Resize procedure: add resize procedure here
Performance measures (RC5) Description: add measures description here
Formulation: add math formulation here
Limitations: add measure limitations here
Robot validation Robot model: add robot model name here
End-effector: add end-effector here
Experiment description: add experiment description here
@@
Legend: OBJL: object localisation; FUNC: functional classification; FUNS: functional segmentation; EPE: end-effector pose estimation; EIS: end-effector interaction synthesis; RC: reproducibility challenge.
@@
Notes: *data and weights of the trained model are recommended to be placed in a repository that favours long-term persistence and accessibility.
Figure 12: Visualization of the main challenges tackled by Affordance Sheet sections. KEYS – RC: reproducibility challenge, RoV: robot validation
The redefinition of the visual affordance problem formulation (see Section II) can also result in experimental validations that ignore datasets and benchmarks of partially overlapping subtasks. For example, works on affordance grounding [11, 90, 93, 95] do not compare the performance of their proposed methods with that of methods designed specifically for affordance segmentation [62, 78, 84, 74], even if the problem formulation is similar [11, 62]. Methods for affordance segmentation output an affordance mask (binary mask) for each action in a predefined set of classes, whereas methods for affordance grounding output a confidence map describing where an action known a priori can take place in the image. Despite these differences, comparing methods for both affordance grounding and affordance detection and segmentation can explain if using action as input (affordance grounding) to a model provides any advantage.
VI-B In support of reproducibility: Affordance Sheets
To promote reproducibility in affordance prediction, we propose the Affordance Sheet, an organised collection of good practices that can facilitate fair comparisons and the development of new solutions (see Table X). Model cards [30] were previously introduced to improve the transparency of methods and to raise awareness about limitations of methods, by describing the characteristics of the method, the details of the experimental setup, the applications or conditions leading to underperformance. Our Affordance Sheet includes Model Cards and adds sections that complement the released information. Fig. 12 shows how each section addresses each reproducibility challenge.
The first section requires to identify which problems the affordance model tackles, helping researchers and users understand what are the competing methods and assess the performance of solutions under the same inputs and conditions. This section also implicitly highlights the considered parts of the overall framework (see Section II-B). When proposing a new problem partially overlapping with another one, previous models can be used or adapted to validate the method. For example, selecting the channel of an affordance segmentation output based on the action considered by the affordance grounding method enables the comparison between methods for affordance segmentation and methods for affordance grounding. To compare the grounding and segmentation outputs, the grounding confidence map can be converted to a binary mask via thresholding; alternatively, the segmentation map can be converted to a confidence map by taking the segmentation output before the argmax operation or by using Gaussian blur.
The second section of the Affordance Sheet describes the datasets (RC1) used by the proposed solutions, aiming to detail their characteristics, share the link where other researchers can find the data, and the license informing about the potential uses of the data. We recommend next benchmarks to also release a detailed description on how to use and visualize data so that researchers can get acquainted with the format. Moreover, we recommend that future benchmarks evaluate models under different conditions, such as generalization to different object instances, different object categories, different object poses, background, and clutter. Benchmarks of models for tasks different from visual affordance prediction, such as COCO for object detection and instance segmentation [160], and CORSMAL for object mass estimation [137], release only the training and validation sets while keeping a private testing set to not bias the designer of the architecture [137, 160]. The release of the testing set can lead researchers to make changes aimed at improving performance scores rather than formulating a contribution that advances the field.
The third section highlights the model characteristics (RC2, RC3) integrating information in the model card (if available). Providing model cards [30], along with the model’s implementation and trained weights, helps detail the description of models supporting other researchers to build upon. When not available, we encourage the reimplementation and retraining of the models as a contribution for the community (e.g. see a previous work that reimplemented, retrained, and release models for affordance detection and segmentation to cope with the lack of previously public available models [4, 20]). As also recommended for datasets, we also encourage providing a link to the trained model’s weights and a license detailing the allowed uses. Without a license, the code is automatically protected by copyright, hence other researchers can not directly use the method implementation to reproduce results or as baseline, as for some previous works [17, 86, 91, 94, 97].
By providing the details the experimental setup used to train and evaluate methods (RC4), the fourth section of the Affordance Sheet is fundamental to correctly use previous methods and develop a solution under the same conditions. Setup conditions include pre-processing and post-processing information such as data splits, resize procedures and data normalisation, and hyper-parameters choice. The lack of these details can result in models with significantly different parameters, and hence leading to unfair comparisons with previous works.
The fifth section of the Affordance Sheet focuses on the performance measures (RC5), the criteria used to validate and compare methods with previous solutions. Providing a stand-alone toolkit implementing the performance measures ensures the replicability of the results across different works while including new methods. For visual affordance prediction, we recommend evaluating the performance of models using more than one measure as this practice provides a more comprehensive analysis while identifying different aspects and limitations of the models. For example, in affordance segmentation, precision focuses on how many of the predicted pixels have the correct class and recall emphasizes how many of the annotated pixels are correctly predicted. Therefore, computing more than one score (and avoiding using a single score aggregating multiple performance measures) reduces the risk of drawing misleading conclusions that are based only on partial results.
The last section describes the validation of the method through a robotic setup. In previous works, few of the methods were validated using a robotic platform [10, 12, 83, 92, 114]. Unlike previous sections of the Affordance Sheet, the robot validation depends on the availability of a robot. When a robot experiment can be performed, we recommend reporting the characteristics of the setup, the robot and end-effector models, and the description of the experiment in terms of object and conditions. This transparent reporting allows researchers to assess methods using a common platform.
VI-C Future direction: benchmarking visual affordance
Reproducibility and advancements in the design of novel solutions has been facilitated by available datasets, benchmarks and competitions in various computer vision tasks, such ImageNet for object recognition [161], COCO for object detection and instance segmentation [160], and BOP for object 6D pose estimation [162]. However, benchmarks for visual affordance predictions, supported by workshops and competitions to favour reproducibility and fair comparisons, are not available. The connection to a robotic task makes benchmarking and reproducibility even more challenging. Nevertheless, solutions based on our generic framework and novel methods can be designed for robotic grasping and manipulation tasks [163], such as in-hand manipulation [164], picking in clutter [165], and human-to-robot object handovers [18], whose benchmarking protocols and competitions are available. Benchmarking protocol specific to visual affordance used in a robotic task can be also designed and included in existing competitions to further promote reproducibility and engagement of both robotic and vision communities.
VII Conclusion
We proposed a unified, comprehensive formulation for visual affordance prediction. Our formulation relates tasks, such as affordance segmentation, affordance detection, and grasping detection, that were defined and tackled independently in the previous literature. Using our new formulation, we link visual affordance prediction with various physical properties of an object and the context in which a robot is operating, such as the specific action to interact with the object with respect to a higher-level task to perform. Additionally, we designed a generic framework relating visual affordance prediction, physical property estimation, and robot actuation. As a case study, we compared the performance of methods estimating the mass of manipulated objects and we discussed the challenges in accurately estimating this property hence relating it to visual affordance for the proposed framework. Although current methods do not yet generalise to unseen objects, the estimation of physical properties can help bridging the gap between robot sensors and actuators. We also introduced the Affordance Sheet, a tool to support reproducibility and fair comparisons of learning-based models for visual affordance prediction. Finally, we discussed future research directions, datasets and benchmarks.

### v2

@@ Appendix B Affordance sheets
TABLE XI: Affordance Sheet, inspired by Model Cards [23], to favour transparency and reproducibility of works for visual affordance predictions conditioned on robotic tasks. Example filled with Multi-FinGAN [100] details.
Multi-FinGAN
Affordance task OBJL FUNC FUNS EPE EIS
Datasets
Name: -
Record link*: https://github.com/aalto-intelligent-robotics/Multi-FinGAN/blob/main/data/download_train_data.sh
Licence: -
Proposed method
(RC2, RC3)
Record link*: https://drive.google.com/file/d/19462M8s3tEXe_1_riHuvQegLxzdX-kl2/view
Code link: https://github.com/aalto-intelligent-robotics/Multi-FinGAN
Model card:
Licence: MIT
Experimental setup
Data splits:
Set Images
Training 3000
Validation -
Testing -
Hyperparameters:
Name Value
batch size 100
learning rate 0.0001
schedule linear after 400 epochs
patience -
optmizer Adam
momentum default
weight decay default
resize -
flip -
Resize procedure: Object centric crops resized to $256\times 256$
Performance measures
Description: Interpenetration: amount of voxels in common between object and end-effector.
Formulation: -
Limitations: Interpenetration does not take into account the predicted pose feasibility.
Robot validation Robot model: Franka Emika Panda
End-effector: Barrett hand
Experiment: Intel RealSense D435 camera looking at the scene at 45 degree viewpoint. The model generates 20 grasps per object and then intersection and quality metric of each grasp are computed. The first physically reachable grasp with lowest intersection and highest quality metric is executed on the real robot. The robot needs to grasp the object and, without dropping it, move to the start position and rotate the hand ±90° around the last joint (success). If the object was dropped during the manipulation, the grasp is considered unsuccessful.
@@
Legend: OBJL: object localisation; FUNC: functional classification; FUNS: functional segmentation; EPE: hand pose estimation; EIS: hand interaction synthesis; RC: reproducibility challenge; ’-’: information not available. Notes: *data and weights of the trained model are recommended to be placed in a repository that favours long-term persistence and accessibility.
TABLE XII: Affordance Sheet, inspired by Model Cards [23], to favour transparency and reproducibility of works for visual affordance predictions conditioned on robotic tasks. Example filled with VLTP [38] details.
VLTP
Affordance task OBJL FUNC FUNS EPE EIS
Datasets
Name: RIO
Record link*: https://drive.google.com/drive/folders/1IAvh8tBGS3WWgV4SbVoqhwCkmyoSFffh
Licence: -
Proposed method
(RC2, RC3)
Record link*: -
Code link: https://github.com/HanningChen/VLTP/tree/main
Model card:
Licence: Apache 2.0
Experimental setup
Data splits:
Set Images
Training 27,696
Validation -
Testing 17,218
Hyperparameters:
Name Value
batch size -
learning rate -
schedule -
patience -
optmizer -
momentum -
weight decay -
resize -
flip -
Resize procedure: -
Performance measures
Description: Mean Intersection over Union (mIoU) evaluates how well a model’s predicted segmentation aligns with the ground truth segmentation by calculating the overlap between the predicted and actual regions, averaged for the selected classes.
Formulation: $\frac{1}{O}\sum_{i=1}^{O}\frac{TP_{i}}{FP_{i}+FN_{i}+TP_{i}}$
Limitations: mIoU does not take into account the similarity in shape between the predicted and annotated segmentation mask
Robot validation Robot model: -
End-effector: -
Experiment: -
@@
Legend: OBJL: object localisation; FUNC: functional classification; FUNS: functional segmentation; EPE: hand pose estimation; EIS: hand interaction synthesis; RC: reproducibility challenge; ’-’: information not available. Notes: *data and weights of the trained model are recommended to be placed in a repository that favours long-term persistence and accessibility.
TABLE XIII: Affordance Sheet, inspired by Model Cards [23], to favour transparency and reproducibility of works for visual affordance predictions conditioned on robotic tasks. Example filled with EgoTopo [43] details.
EgoTopo
Affordance task OBJL FUNC FUNS EPE EIS
Datasets
Name: EPIC-Kitchens
Record link*: https://data.bris.ac.uk/data/dataset/2g1n6qdydwa9u22shpxqzp0t8m
Licence: CC-BY-NC 4.0
Proposed method
(RC2, RC3)
Record link*: https://dl.fbaipublicfiles.com/ego-topo/anticipation/pretrained.zip
Code link: https://github.com/facebookresearch/ego-topo
Model card:
Licence: CC-BY-NC 4.0
Experimental setup
Data splits:
Set Images
Training -
Validation -
Testing 1,155
Hyperparameters:
Name Value
epochs 20
batch size 256
learning rate 0.0001
schedule 0.1x after 15 epochs
patience -
optmizer Adam
momentum -
weight decay 0.000001
resize -
flip -
Resize procedure: -
Performance measures
Description: Mean average precision (mAP) over all afforded interactions.
Formulation: -
Limitations: mAP score weights equally the classes, regardless of their frequency.
Robot validation Robot model: -
End-effector: -
Experiment: -
Legend: OBJL: object localisation; FUNC: functional classification; FUNS: functional segmentation; EPE: hand pose estimation; EIS: hand interaction synthesis; RC: reproducibility challenge; ’-’: information not available. Notes: *data and weights of the trained model are recommended to be placed in a repository that favours long-term persistence and accessibility.